OpenStudy (yacoub1993):

What is the simplest form of the expression? ∛45 + ∛2058 - ∛750

OpenStudy (anonymous):

\[\sqrt[3]{45}\] cannot be "simplified" in any way

OpenStudy (yacoub1993):

A. 9∛6 B. 2.8∛6 C. 4∛6 D. 14∛6

OpenStudy (anonymous):

\[2058=6\times 7^3\] so \[\sqrt[3]{2058}=7\sqrt[3]{6}\]

OpenStudy (anonymous):

\[750=6\times 5^3\] so \[\sqrt[3]{750}=5\sqrt[3]{6}\]

OpenStudy (anonymous):

but we are stuck with \(\sqrt[3]{45}\)

OpenStudy (anonymous):

either there is a typo on the question, or it is a straight up mistake made by whoever wrote it

OpenStudy (yacoub1993):

http://prntscr.com/34ptn5

OpenStudy (yacoub1993):

@a1234 need ur help please

OpenStudy (a1234):

I'm sorry, I don't know.